The velocity of a train changes from 0 to 60km/hr in 20 seconds. Assuming that it is


uniformly accelerated, what is the acceleration in m/s2?

1
Expert's answer
2022-05-22T14:37:26-0400
"a=\frac{\Delta v}{\Delta t}=\frac{60\div 3.6-0}{20}=0.83\:\rm m/s^2"